# 3D Anderson localization of light in disordered systems of dielectric particles

### v1

@@ 3 Results and discussion
Transverse localization in a dense structure may be accompanied by the changes in the time dependence of transmission, analogous to those observed for dense packings of PEC spheres in [34]. Here we perform a related analysis. We simulate time-resolved transmission by layers $T(t)$ varying their turbidity, i.e. the volume fraction and the refractive index. At first, we consider layers with perturbed side boundaries and with dimensions of $kR=32$ and $kL=14.5$. The volume fraction $\rho$ is varied between 0.17 and 0.5.
@@
Non-exponential decay can be formally represented by means of a time-dependent diffusion coefficient $D(t)$. At the localized state, it decreases as $t^{-1}$ according to the solution of the diffusion equation for the case of $kl^{*}<1$ in time domain [39]. Figure 3d shows time resolved diffusion coefficients $D(t)$ deduced from the computed $T(t)$ by local exponential fitting. One can see a transition from small variation of $D(t)$ in sparse structures to a steep function at $\rho>0.44$ and times $t>15t_{D}$.
Evidently we would have to extend the observation time limit in order to study the evolution of $T(t)$ at longer times and to find out if $D(t)$ can reach the predicted $t^{-1}$ dependence. This could be achieved by increasing sample thickness and diameter roughly by a factor of two or more. This would correspond to $kL\gtrsim 30$ and a number of particles $N>10^{5}$ which makes the problem computationally extremely expensive, even for modern large scale HPC clusters. Another way to hinder energy decay in time is to create conditions for a whispering-gallery effect at the side boundary of a cylindric layer. For this purpose we did computations for layers with smooth boundaries and a particle size of $kr=1$. The results for $T(t)$ and $D(t)$ dependencies at different volume fractions from $0.1$ to $0.5$ are shown in Figures 4ab. The dimensions of the layers are $kR=26$ and $kL=12$. A densely packed sample with $\rho=0.44$ is shown in the inset in Figure 4a. Figures 4cd represent snapshots of the electric field distributions ($E_{x}$ field component) in the horizontal XY-plane at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers. The WG waves, excited by the incident pulse due to the smooth boundaries of samples with cylindric symmetry, are coupled to the side boundary due to total internal reflection and allow us reaching $t\approx 40t_{D}$ for relatively small samples. At the same time the field distributions remain random in the bulk. Despite the WG effect, present in both cases, the structures show different transmission properties. The $T(t)$ plot in Figure 4a clearly shows a transition from purely exponential decay to localization at large volume fractions and times $t>20t_{D}$. This occurs at $\rho\gtrsim 0.3$. The WG effect is observed in all the layers above and below this value, even in the sparse structures for which the the condition $l^{*}\lesssim\lambda$ holds. This proves that it is not the WG mode itself but the particle packing density which is causing the non-exponential behaviour of $T(t)$. The diffusion coefficient $D(t)$ at times $t\gtrsim 20t_{D}$ obtained for the dense structures with $\rho\gtrsim 0.44$ follows a $t^{-1}$ fit in Figure 4b demonstrating that we have reached the AL regime. We note that this result is consistent, in terms of time and the volume fraction threshold, with that obtained for layers of PEC spheres in [34].
Figure 5: (a) Normalized transmission of a short pulse through a layer with a volume fraction $\rho=0.44$ and a smooth boundary as a function of time $T(t)$ at different refractive indices. (b) Time-dependent diffusion coefficient $D(t)$ as obtained by a local exponential fitting of $T(t)$. A localization signature $D(t)\sim t^{-1}$ emerges at times $t>20t_{D}$ and $n=3.0$.
@@
Figure 6: A snapshot of the near field $|E|^{2}$ excited by the propagation of a short pulse through a dense layer with $\rho=0.48$ and $n=3.0$ shown in Fig. 1a along $Z$ axis at times $t=2t_{D}$ (a) and $t=20t_{D}$ (b). A fraction of light is trapped by the disordered air-dielectric interface through evanescent coupling. This results in a constant flow along complex, highly disordered paths which slows down the energy decay through the layer boundaries. An animation showing the time dynamics of the near field can be found in the Supplementary material A.1.
Our results imply a mechanism that slows down the energy decay at dense packing. Apparently, it should be connected with the percolation-like light transport in a thick, dense layer we analyzed above. Visualization of the field dynamics excited by a short pulse yields a clue to this mechanism. One can conventionally split up the propagating field into two components. The first one is the wave packet of the incident pulse that quickly diffuses through a layer. The second one is created by the energy leaked from the packet through the evanescent field-coupling. Being coupled once, this component remains trapped by the air-dielectric interface, which, together with the free-space cavities, represents a geometrically complex 3D structure. Despite little free space at dense packing, such a structure possesses a long-range connectivity, i.e. arbitrary points in a void can be connected with another point in an arbitrary free-space location. The paths connecting such arbitrary points can be much longer than the sample dimensions.
@@
In Fig. 6, one can see snapshots of the near-field intensities inside the thick layer shown in Fig. 1a at a time right after the pulse passage $t=2t_{D}$ and at the longer time $t=20t_{D}$. An animated field evolution for a period of time after $2t_{D}$ can be found in the Supplementary material A.1. The trapped field uniformly fills the sample volume and propagates along complex paths, constantly changing direction and forming hotspots in the voids. The paths can be not necessarily closed loops, but they are apparently sufficiently long. The field decay occurs only when the field reaches the sample boundaries. Thus, the energy in the interior needs substantial time in order to reach the boundary, much longer than the time of propagation along a straight line or a sequence of straight lines between scattering events in a sparse medium. The character of such an internal field evolution does not significantly change at longer times, and the decay of hotspots is not monotonous in each location. They demonstrate different behaviour in such a disordered system: those that are excited directly by the transmitted pulse decay while others arise in new locations. The decay rate is also different for different hotspots, i.e. they may have different life times. Some are decaying within $\approx 5$ oscillation periods while others are much more stable. Non-monotonous decay and enhancement with time may also take place in the same location, which proves a high geometrical complexity of the paths. If the energy of a path is leaked at the sample boundary, it is lost in the total decay. A leakage in the interior transfers energy to new paths with similar spatial scales and degrees of complexity, forming a complex percolation graph. Thus, the entire process appears to be a sophisticated spatiotemporal 3D flow of energy and should be a subject for further studies. In particular, such a mechanism can be relevant for low-absorbing materials as it involves field-interface interaction and propagation in free space. The role of the refractive index also becomes clear in this situation. It determines the balance between the energy transmitted by particles and chains of particles, enhancing a diffusive process, and the energy flowing through a random system of connected cavities with long path lengths that participates in localization.
4 Conclusions
We observe clear signatures of AL in the results of the full wave numerical simulations of light transmission by 3D dielectric particulate layers with high degrees of structural disorder. A transition from diffusive transport to localization occurs at approaching the Ioffe-Regel criterion for the mean free path length $kl^{*}\lesssim 1$ in densely packed systems and at the refractive index of $n=3.0$, which is consistent with the predictions of theoretical models. Simulation of the focused beam propagation in a dense layer demonstrates an absence of transverse spreading and anomalous conservation of the beam’s linear polarization state in contrast to the classical light scattering in a sparse analog structure. Time-resolved transmission of a short pulse in layers with similar densities and constituent particle sizes exhibits a non-exponential energy decay already at times $t>15t_{D}$. For observations at $t\gg t_{D}$ larger structures must be considered which makes the problem computationally very expensive even using HPC. Alternatively, we extend the observation time retaining sample cylindric symmetry. This allows excitation of whispering gallery modes that hinder field decay from a sample. Therewith we found that the time-resolved diffusion coefficient varies in agreement with the self-consistent theory of localization at times $t>20t_{D}$ and volume fractions $\rho\gtrsim 0.44$. Thus, Anderson localization of light is possible in dense highly disordered 3D dielectric structures which motivates a continued experimental search of the AL characteristics in uncorrelated disordered media, with realistic material refractive indices not exceeding $n=3.0$.

### v4

@@ Abstract
Transverse localization in a dense structure may be accompanied by the changes in the time dependence of transmission, analogous to those observed for dense packings of metallic spheres in Ref. [22]. Here we perform a related analysis. We simulate time-resolved transmission by layers $T(t)$ varying their turbidity, i.e. the volume fraction and the refractive index. At first, we consider layers with perturbed side boundaries and with dimensions of $X_{R}=32$ and $X_{L}=14.5$. The volume fraction $\rho$ is varied between 0.17 and 0.5. An example of a dense layer with $\rho=0.44$ and its cross-section are shown in Figures 2ab.
AL is expected for materials with high refractive indices. E.g., TiO2 powders are reported as those where transverse localization and anomalous transmission can be observed [31]. Therefore, we first fix the refractive index at $n=3.0$ to study the role of the volume fraction factor. The computed transmissions $T(t)$ for different samples are shown in Fig 2c. The layers are illuminated by a short Gaussian pulse. The arrival of the transmitted pulse intensity peak at the detector plane is recorded as time $t_{a}$ for each case. The horizontal axis represents normalized time $t/\penalty 50t_{a}$. The maximum observation time is limited to $\approx 25t_{a}$ as the samples are relatively thin and the measured intensity quickly reaches the numerical accuracy of double-precision numbers. However, we observe a non-exponential dependence of $T(t)$ for larger volume fractions near this limit. With increasing turbidity at larger $\rho$, we approach the $kl^{*}\lesssim 1$ condition obtaining not only a decreasing slope of $T(t)$ but a noticeable deviation of transmission from exponent at longer times.
Non-exponential decay can be formally represented by means of a time-dependent diffusion coefficient $D(t)$. At the localized state, it decreases as $t^{-1}$ according to the solution of the diffusion equation for the case of $kl^{*}\lesssim 1$ in time domain [25, 22]. Figure 2d shows time-resolved diffusion coefficients $D(t)$ deduced from the computed $T(t)$ by local exponential fitting. One can see a transition from small variation of $D(t)$ in sparse structures to a steep function at $\rho>0.44$ and times $t>15t_{a}$.
Evidently we would have to extend the observation time limit in order to study the evolution of $T(t)$ at longer times and to find out if $D(t)$ can reach the predicted $t^{-1}$ dependence. This could be achieved by increasing sample dimensions roughly by a factor of two or more and, correspondingly the number of particles to $N\gtrsim 10^{5}$, which makes the problem computationally extremely expensive, even for modern large scale HPC clusters. Another way to hinder energy decay in time is to create conditions for a whispering-gallery effect at the side boundary of a cylindric layer. For this purpose we did computations for layers with smooth side boundaries. The results for $T(t)$ and $D(t)$ dependencies at different volume fractions from $0.1$ to $0.5$ are shown in Figures 3ab. The dimensions of the layers are $X_{R}=26$ and $X_{L}=12$. A densely packed sample with $\rho=0.44$ is shown in the inset in Figure 3a. Figures 3cd represent snapshots of the electric field distributions ($E_{x}$ component) in the horizontal XY-plane at time $t=25t_{a}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers. The WG waves, excited by the incident pulse, allow us reaching $t\approx 40t_{a}$ for the samples with large volume fractions and relatively small thickness. At the same time the field distributions remain random in the bulk. Despite the WG effect, present in all cases, including the sparse structures, they show different transmission properties. The $T(t)$ plot in Figure 3a clearly demonstrates a transition from exponential decay to localization at large volume fractions $\rho\gtrsim 0.3$ and times $t>20t_{a}$.
The diffusion coefficient $D(t)$ at times $t\gtrsim 20t_{a}$ obtained for the dense structures with $\rho\geq 0.44$ follows a $t^{-1}$ fit in Figure 3b demonstrating that we have reached the AL regime. This result is consistent, in terms of time and the volume fraction threshold, with that obtained for layers of metallic spheres in [22]. Interestingly, a WG mode is established at times $t\approx 10t_{a}$, prior to the emergence of the $D(t)\sim t^{-1}$ regime. This proves that it is not the WG mode itself but the particle packing density which is causing the non-exponential behavior of $T(t)$. This is also confirmed by the results of simulations with periodic boundary conditions presented in the Supplementary material III.
Figure 3: Simulation results for layers with different volume fractions and smooth side boundaries. The refractive index of the material is $n=3.0$. (a) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.1$ to 0.5 as a function of time $T(t)$. (b) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Dense layers demonstrate a $t^{-1}$ (dashed lines) dependence at times $t>20t_{a}$. (c) and (d), Snapshots of the internal electric field (XY-cross-sections, $E_{x}$ component) distributions showing a whispering-gallery wave propagating along the boundary at time $t=25t_{a}$ for a moderately sparse ($\rho=0.27$) and a dense ($\rho=0.44$) layers. (e) Normalized spectra of transmitted light as functions of the dimensionless size parameter $X_{r}$ (inverse wavelength expressed relative to the particles size) for sparse $\rho=0.17$ and dense $\rho=0.44$ layers. Dashed line corresponds to a Lorentzian fit for $\rho=0.44$. The inset shows the Thouless conductance distribution $P(\textrm{ln}~g_{Th})$ for $\rho=0.44$ that has a mean value $\mu=-1.25$ and a standard deviation $\sigma=0.75$.
To make the characterization of the localized transport more complete we compute transmission spectra for the sparse ($\rho=0.17$) and dense ($\rho=0.44$) samples (Figure 3e). For the horizontal axis a dimensionless size parameter $X_{r}$ is used, which is equivalent to the inverse wavelength relative to the particle size. The $X_{r}$ range in the plot spans approximately the full width at half maximum of the incident pulse. The $\rho=0.17$ spectrum is represented by broad, overlapping, maxima. The spectrum for $\rho=0.44$ consists of non-overlapping sharp peaks resulting from the states spatially confined in the medium [19, 26, 27, 28, 29]. Dashed line shows a spectral fit obtained with an ensemble of Lorentzian oscillators. The separations of peaks are larger than their widths, which indicates that the Thouless criterion for AL is satisfied. With this, a Thouless conductance $g_{Th}$ can be defined as the ratio $g_{Th}=\delta\omega/\Delta\omega$, where $\delta\omega$ is the average width of two adjacent modes in the spectrum and $\Delta\omega$ is their spectral separation. Thus, $g_{Th}<1$ corresponds to a localized propagation. In the case of $\rho=0.44$, where the signatures of localization emerge, we estimate it as an average over all visible modes as $g_{Th}>\approx 0.36$. The inset in Figure 3e shows the distrubution $P(\textrm{ln}~g_{Th})$. It approximately follows a Gauss law according to the theoretical predictions [28].
Figure 4: (a) Normalized transmission of a short pulse through a layer with a volume fraction $\rho=0.44$ and a smooth boundary as a function of time $T(t)$ at different refractive indices. (b) Diffusion coefficient $D(t)$ as obtained by a local exponential fitting of $T(t)$. Dashed line shows a $t^{-1}$ dependence.
Another free parameter of the medium that is important for light transport is the refractive index of the material. To illustrate the transition between diffuse and localized regimes from this point of view, we do similar computations of $T(t)$ for the dense sample with $\rho=0.44$ and refractive indices ranging from n=n=2.0 to 3.0. Figure 4 shows that $n$ must be larger than 2.5 for a non-exponential decay, and further increase of $n$ enhances this effect at longer times. This is also consistent with theoretical estimations and experimental measurements for materials with different $n$ that suggest $n\gtrsim 2$ [9, 31]. Thus, the refractive index contrast is another critical factor for AL in dielectric structures that can impede or enhance the localized transport. It turns out that $n$ should not be necessarily extremely large and a continued experimental search for AL in ’white paint’ powders like TiO2 is not futile.
Our results suggest a mechanism that slows energy decay under dense packing conditions, likely linked to percolation-like transport. We provide our intepretation of the observed transition in the Supplementary material III , based on the analysis of the near-field dynamics. After a short pulse passes through the system, a decaying field component becomes trapped at the air–dielectric interface, exhibiting a complex three-dimensional spatiotemporal energy flow. In a strongly disordered medium with long-range connectivity, propagation paths can be intricate and significantly exceed the physical thickness of the sample. Coherent transport along such extended paths, emerging from an initially random near-field distribution, leads over time to the formation of stable, disorder-induced spatially localized modes. These modes have sufficiently long lifetimes and are responsible for the sharp peaks observed in the transmission spectrum.
We note, that this highly complex field propagation cannot be adequately described by multiple light scattering in a cloud of identical point scatterers. In the last case, a vector treatment shows that near-field coupling and interference between the longitudinal and transverse components suppress AL [40, 39]. However, it has also been shown that the amplitude of the longitudinal component, specific to point scatterers, non-monotonously depends on size for spheres and conditions can even be created to minimize it [52]. In contrast to point-scatterers and spheres, irregular particles produce inherently non-symmetric dipolar near fields with random amplitudes determined by the particle orientation relative to the polarization of the incident wave. Consequently, the transverse and longitudinal components acquire random amplitudes, and their interference does not necessarily prevent AL. Nevertheless, the problem of the longitudinal field remains an important question worth a separate consideration. It would be particularly interesting to study its role in systems with different types of constituents and finely tunable spatial correlations. This can be very useful for understanding the dynamics of AL in 3D, especially since the local neighbourhood geometry is an important factor for an initial near field distribution and its subsequent evolution.
In conclusion, our full-wave numerical simulations of light transmission through 3D disordered dielectric particulate layers reveal clear signatures of AL. We observe a transition from diffusive transport to localization as the system approaches the Ioffe–Regel condition at high volume fractions ($\rho\gtrsim 0.44$) and a refractive index of $n=3.0$, consistent with theoretical predictions. Simulation of focused beam propagation in a thick, dense layer demonstrates an absence of its transverse spreading. Time-resolved transmission of a short pulse in layers with similar volume fractions and particle sizes exhibits a non-exponential energy decay with the time-dependent diffusion coefficient varying in accordance with the self-consistent theory of localization. The transmission spectrum satisfies the Thouless criterion for AL in this case. Near-field visualizations reveal a clear dynamic transition from diffusive propagation to the formation of spatially localized modes. These findings demonstrate that Anderson localization of light can occur in highly disordered 3D dielectric structures, motivating further experimental efforts to identify its hallmarks in realistic disordered media with refractive indices not exceeding n = 3.0.
Acknowledgements.
The authors gratefully acknowledge the computing time granted by the Paderborn Center for Parallel Computing (PC2).
Supplementary material
I Numerical model
We performed numerical simulations using a self-developed light scattering code based on the open source DGTD Maxwell solver MIDG published by Tim Warburton [53]. The CST Microwave Studio software package based on the Time Domain Finite Integration Technique was used for a qualitative benchmarking the DGTD simulation results for simplified setups. In our code, we implemented plane-wave and focused beam lights sources, boundary conditions and material description. Here, we use continuous and Gaussian pulses. A plane wave source is realized using the total field/scattered field technique. For simulations with a focused beam we implemented the plane wave superposition method [54]. An open computational domain is pretended with the PML absorbing boundary condition. For the time-resolved transmission measurements we illuminated target layers with a linearly polarized short plane wave pulse. The intensity of the transmitted signal is recorded, correspondingly, in a monitor plane behind the layer. The intensity curves in all plots are normalized by the peak intensity of the transmitted pulse. For generation of the particulate samples with different dimensions and controlled filling fractions we use the Bullet Physics engine [50]. This is an open-source C++ library that can be also used as an add-on in the Blender software. It allows simulation of dynamics and collision of multiple arbitrary 3D shapes in time domain. Thus, simulating natural powder mechanics samples with realistic topologies can be created. For sparse distributions of the pre-generated random irregular shapes we simulated their free fall on a substrate in a closed cylindrical volume. The sets of particles are created with a self-developed shape generator based on the Gaussian random field approach [49]. 3D models of the generated layers were then used for tetrahedral mesh generation. The spatial resolution in the regions with dielectric material with $n=3.0$ was $\sim 15$ tetrahedral cells per central wavelength of the incident pulse and the nodal expansion order in the DGTD numerical scheme was $N=3$. All the DGTD simulations were done on the HPC cluster Noctua 2 of Paderborn University using up to 30 cluster nodes per simulation.
II Short pulse transmission by layer samples with reduced dimensions
We take a cylindrically symmetric sample with $X_{R}=26$, $X_{L}=12$ and $\rho=0.44$ as a reference and consider reduced structures with the same bulk volime fraction and thicknesses $X_{L}=6$ and 3 and an elementary bi-layer structure. The refractive index is fixed at $n=3.0$. The results of the simulated transmission $T(t)$ for four layers are presented in Figure S1a. A bi-layer shows exponential decay even though a WG mode is excited. The case of $X_{L}=3$ is equivalent to a doubling the thickness of the bi-layer sample. This appears to be enough to make the $T(t)$ curve non-exponential. In all four cases a WG mode is excited allowing us reaching times $t>20t_{a}$. Further thickness increasing leads to stronger deviations from exponent and more rapid reduction of the diffusion coefficient $D(t)$ (Figure S1b). However, only with $X_{L}=12$ we obtain a qualitative change at longer times and reach the $t^{-1}$ fit.
The whispering-gallery field pattern that we observe in all samples with cylindrical symmetry at longer times is, in fact, a complex 3D oscillation both along the round side boundary and in the vertical direction (along $Z$) between the upper and bottom sides of the cylinder. This is indeed a whispering-gallery process caused by cylindrical symmetry but not entirely a classic WG-mode that is well known for, e.g., solid micro-discs. There is also a constant flow of energy in the central part of the sample volume that interacts with the WG waves which is not typical for solid structures with circular symmetries. The random flow in the central part and in the bulk is a result of constant evanescent field coupling to a random network of dielectric interfaces feeded by the WG-mode. Such a near-field propagation process deserves a separate study as, to our knowledge, phenomena of this kind have not been described in literature yet.
Figure S1: (a) Normalized transmission of a short pulse through layers with cylindric symmetry and different thickness $X_{L}$ as a function of time $T(t)$. The diameter and volume fraction are $X_{R}=26$ and $\rho=0.44$, correspondingly, and the refractive index is $n=3.0$. The energy decay from a simple double layer of particles is exponential while thicker layers demonstrate increasingly non-exponential transmission. (b) Diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. $D(t)$ evolves from a constant to a $t^{-1}$ fit (dashed line) with increasing thickness. (c) Vertical cross-sections of the samples with different thickness.
Figure S2: Normalized transmission of a short pulse through a small layer of 2000 particles packed with volume fraction $\rho=0.5$ (inset) and with dimensions $X_{R}=15$ and $X_{L}=3$ as a function of time $T(t)$ at different refractive indices. Simulations are done with the CST Microwave Studio software package.
For the computations with the CST FIT solver we take a cylindric sample with thickness $X_{L}=3$ (a four-particle thick layer) and reduced diameter $X_{R}=15$. Due to parallelization restrictions of CST MWS we need to simplify the target system. With the particle size $X_{r}=1$ and the volume fraction $\rho=0.5$ the number of particles becomes relatively small in this case, N=2000. With fixed sample parameters we vary the material refractive index $n$ from 2.0 to 3.0. For the CST computations we use a polyhedral description of the model sample and, then, a hexahedral mesh is generated by the software internally. The electromagnetic problem is solved in time domain simulating propagation of the short plane-wave pulse, similarly to our DGTD method setup.
The result in Figure S2 is qualitatively similar to the $T(t)$ refractive index dependence in Figure 4a obtained for a larger system with the DGTD method.
III Analysis of the near field dynamics
In this section we analyze light transport in a thick layer with a large number of particles, shown in Figure 1a, in more detail and consider a problem of light transmission in layers with reduced dimensions and periodic boundary conditions, in order to interpret our observations of the localization signatures presented in Figure 3.
At first, we visualize the excited near-field distribution in the thick sample and animate its evolution in time. We additionally simulate propagation of a short focused pulse in the same layer sample with the same model parameters. Video III represents an $XZ$-cross-section of the computational domain with the near-field intensity distribution. It shows a CA-like behavior of the field hotspots in the sub-wavelength voids between the closely neighboring particles. We note that the wave packet is partially preserved along the propagation path. Its effective wavelength becomes smaller, $\lambda_{eff}\approx 0.6\lambda_{inc}$, increasing the effective size of particles to $X_{r,eff}\approx 1.8$. Nevertheless, the condition $kl^{*}\lesssim 1$ or, at least, $kl^{*}\sim 1$ is preserved for most of the voids.
{video}
@@ III Analysis of the near field dynamics
Additional characterization can be obtained from the analysis of the polarization state of the near field. Generally, diffusive propagation in a disordered medium depolarizes incident radiation after many scattering events that are strongly stochastic in a disordered multi-particle system. A steady-state distribution of the electric field component $E_{x}$ in Fig. S3 computed for the same dense layer and same conditions as in Figure 1b shows a counter-intuitive conservation of the original polarization state of the propagating beam at the exit. The $E_{y}$ and $E_{z}$ components are excited along the path and play a role in the field coupling process but they do not propagate directionally. This fact can explain the halt of transverse spreading: the propagating beam ”prefers” cavities where the initial polarization direction can be preserved and, hence, the $\vec{k}$-vector bending is avoided and the direction of propagation is not changed.
Figure S3: Steady-state distributions of the three components of the electric field at the propagation of a continuous focused beam through the dense layer with $\rho=0.48$ and $n=3.0$ shown in Figure 1a.
From the focused short- and continuous-beam simulations we see that the fundamental property of the observed transport mechanism in densely packed layers is concentration of the transported energy in the free-space cavities, while less is refracted inside the dielectric particles. A wave can couple to the neighbor-particle interface while being still coupled to the previous one propagating in a void between two interfaces. On the other hand, these interfaces are disconnected and belong to different scatterers. Thus, the field appears to be constantly coupled and one can say that there are no more classic ”scattering events”. Therefore, the term ”free path length” becomes irrelevant in its classic definition. The condition $kl^{*}\lesssim 1$ implies also an extremely small mean free time $t^{*}$, and the near field, being coupled to a particle, most often reaches its neighbour at time $t\ll T$, where $T$ is the wave oscillation period. We note also, that the Ioffe-Regel criterion should be considered as qualitative for the localization problems [15].
The parameter $kl^{*}$ can be measured only indirectly and the existing methods for such indirect estimations should be, therefore, validated for the case of the vector field propagating in a dense 3D medium. This is, however, beyond the scope of this paper. Thus, we use $kl^{*}$ just to characterise the size of the voids in a particulate medium with respect to the wavelength. We only roughly estimate it geometrically and then relate it to the Ioffe-Regel condition. The estimation follows from a simple consideration that at volume fractions $\rho=0.45–0.5$, which is close to the theoretical maximum for mono-disperse irregular constituents, the characteristic size of voids between particles cannot be larger than the particle size. The particle sizes in our simulations are $X_{r}=1–1.1$ or $X_{r,eff}\approx 1.8$ with respect to the effective wavelength at $\rho=0.44$ and $n=3.0$. Correspondingly, the characteristic free path length cannot be larger than that.
For the further explanation of the light transport process at longer times, it is convenient to split up the propagating field into two components. The first one is the wave packet of the incident pulse that quickly passes through a layer. The second one is created by the energy leaked from the packet through the evanescent field-coupling. A field evolution video (Video III) can better represent the dynamics of the arrested component in a short period of time after $2t_{a}$. Randomization of the propagation direction, required for localization at longer times, is realized through coupling to the random air-dielectric interface. Being coupled once, the delayed component fills the sample volume and remains trapped by the interface, which represents a geometrically complex 3D structure. Despite little free space at dense packing, a particulate layer possesses a long-range connectivity, i.e., an arbitrary point in a void can be connected with another point in an arbitrary free-space location. Thus, field propagates along complex paths, constantly changing direction and forming randomly distributed hotspots in the voids (Figure S4, $t=2t_{a}$).
{video}
@@
Figure S4: Snapshots of the normalized near field $|E|^{2}$ excited by the propagation of a short pulse through a dense layer with $\rho=0.48$ and $n=3.0$ shown in Figure 1a along $Z$ axis at times $t=2t_{a}$ (a) and $t=20t_{a}$ (b).
Figure S5: Cross-sections of the computational domain showing normalized near field $|E|^{2}$ in the sample with cylindric symmetry (Fig. 3a) and $\rho=0.44$ at time $t=25t_{a}$.
@@
The field decay occurs only when the field reaches the sample boundaries. Thus, the energy in the interior needs substantial time in order to reach the boundary, much longer than the time of propagation along a straight line or a sequence of straight lines between scattering events in a sparse medium. The decay of hotspots is not monotonic in each location. They demonstrate different behavior in such a disordered system already at earlier times (Video III): those that are excited directly by the transmitted pulse decay while others arise in new locations. The decay rate is also different for different hotspots, i.e. they may have different life times. Some are decaying within $\approx 5$ oscillation periods while others are much more stable. Non-monotonic decay and enhancement with time may also take place in the same location, which proves a high geometrical complexity of the paths. If the energy of a path is leaked at the sample boundary, it is lost in the total decay. A leakage in the interior transfers energy to new paths with similar spatial scales and degrees of complexity, forming a complex percolation graph. Evanescent coupling plays a crucial role here. Small dielectric particles act as resonant scatterers and the evanescent field tails from their surfaces overlap with neighboring particles. Such overlap is phase-sensitive and provides coherent pathways in a dense system.
The character of such an internal field evolution does not significantly change in time but the hotspot distribution eventually becomes inhomogeneous. This becomes apparent in Figure S4 at $t=20t_{a}$ and it is even more clearly seen in the near-field inside a cylindric sample, for which we observe a non-exponential tail in transmission at longer times. Figure S5 shows $XY$-, XZXZ- and $YZ$-cross-sections of the computational domain with the near-field intensity for the $\rho=0.44$ sample at $t=25t_{a}$. Together with the WG pattern along the boundary, one can see notice clustering of the intensity hotspots in the bulk.
To make sure that such a structuring is preserved in the absence of the WG mode, we did a similar series of simulations for a square layer with periodic boundary conditions. All model parameters, such as gaussian pulse, particle size and refractive index ($n=3.0$) remain the same, except reduced dimensions of a unit cell ($X_{d}=18$ along X and Y axes and $X_{L}=9.6$ along Z). We vary volume fraction $\rho$ from 0.18 to 0.44. Periodic boundary conditions is another way to hinder energy decay from a finite sample. However, one should be careful and avoid phenomena like standing waves and guiding modes, that may appear over long time in an artificially infinite thin layer and influence the $T(t)$ curve.
Figure S6: a. Normalized transmission of a short pulse through periodic layers with different volume fractions (inset: $\rho=0.44$) as a function of time $T(t)$. b. Diffusion coefficient $D(t)$ deduced from the $T(t)$ data. c. Transmission spectra for different volume fractions measured at longer times and at the normalized transmission level of $T\approx 10^{-9}$. d. XYXY-snapshots of the near-field intensity $|E|^{2}$ inside layers with different volume fractions at times corresponding to the normalized transmission $T\approx 10^{-8}$.
{video}